# Quantum property testing in sparse directed graphs

### v1

@@ 4.2 Preparation
Notice that this problem is not a property testing problem, as the outcome is not determined based on the distance between inputs. Nevertheless, it is a valid promise problem and a special case of testing $k$-collision-freeness, that we use to prove a lower bound on the other problems of interest.
To prove a bound on $\mathrm{Collision}$, we will actually relate it to the composition of two more elementary functions, where by composition we mean $(f\circ g)(x)=f(g(x_{1}),\dots,g(x_{n}))$, for appropriate functions $f$ and $g$. Define (i) the threshold function $\mathrm{THR}_{N}^{k}:\{-1,1\}^{N}\to\{-1,1\}$ which is $-1$ if the input bitstring contains at least $k$ many $-1$s, and it is 1 otherwise; and (2) the gap version of $\mathrm{OR}$, that is $\mathrm{GapOR}_{R}^{\gamma}:D_{\mathrm{GapOR}_{R}^{\gamma}}\subset\{-1,1\}^{R}\to\{-1,1\}$ which takes value 1 if the input is $1^{R}$, $-1$ if the input contains at least $\gamma R$ many $-1$s, and is undefined otherwise. We show that the double-promise approximate degree of $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ lower bounds the quantum query complexity of the collision problem.
@@
Notice that $(s_{1},\dots,s_{N})$ is free from $k$-collisions (ignoring collisions of the dummy character 0) if and only if $(T_{1}(s_{1}),\dots,T_{N}(s_{N}))$ is free from $k$-collisions, i.e. new $k$-collisions cannot be created by this transformation (only 2-collisions but we assume $k\geq 3$).
@@
Using Proposition 4.11, instead of $\mathrm{Collision}^{k,\gamma/c}_{N,N}$ we can consider $\mathrm{dCollision}^{k,\gamma}_{N,R}$ (with the appropriate parameters) to show a lower bound. By Proposition 4.12, we can use Lemma 4.7 to relate the query complexity of $\mathrm{dCollision}^{k,\gamma}_{N,R}$ to the double-promise degree of $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ with domain further restricted to $H_{\leq N}^{NR}$. ∎
@@ 4.3 Main lower bound
We first define a construction used in order to compose dual polynomials.
@@
Let $N=\lceil 20(2k)^{k/2}\rceil R$ and $\phi:\{-1,1\}^{R}\to\mathbb{R}$ from Proposition 4.15. Then there exists $\psi:\{-1,1\}^{N}\to\mathbb{R}$ and $\zeta:\{-1,1\}^{NR}\to\mathbb{R}$ such that
@@
As now we focus on the composed function $f$ (and the dual composition $\phi\star\psi$), the domain is not restricted to small Hamming weight inputs anymore.
@@
Then, similarly to the first term, we can upper bound the probability. Now in the last step we use the Chernoff bound, which introduces the constraint $\gamma<\frac{1}{4^{k-1}}$.
@@
By Lemma 4.14 there is a dual polynomial for $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ of pure high degree $\Omega(\sqrt{N^{1-1/k}}/\ln^{2}N)$ that is only supported on $H_{\leq N}^{NR}$. By Theorem 4.8 this means that the double-promise $\delta$-approximate degree of $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ with domain restricted to $H_{\leq N}^{NR}$ is $\Omega(\sqrt{N^{1-1/k}}/\ln^{2}N)$. Using Lemma 4.10 with $c=\lceil 20(2k)^{k/2}\rceil$ we obtain that the bounded-error quantum query complexity of $\mathrm{Collision}_{N,N}^{k,\gamma^{\prime}}$ is $\Omega(\sqrt{N^{1-1/k}}/\ln^{2}N)$ if $\gamma^{\prime}=\gamma/c<1/(4^{k-1}\lceil 20(2k)^{k/2}\rceil)$. This implies the same lower bound on testing $k$-collision-freeness with $\varepsilon=\gamma^{\prime}$ as $\mathrm{Collision}$ is just a more restricted version of the same problem. ∎
@@ 5 Testing $3$-colorability
In this section, we will prove that the problem of testing $3$-colorability in bounded degree graphs remains maximally hard-to-test in the quantum setting. Our lower bound proof will roughly follow the same approach as that of [bogdanov2002lower]. See [bhattacharyya2022property, Section 5.6] also for a reference.
@@
Let us start with the notion of $k$-wise independent string.
@@
In order to prove the above theorem, we will be using the following observation from [apers2022quantum], which states that distinguishing between uniformly random string and $\ell$-wise independent string, for an appropriate integer $\ell$, is hard for quantum algorithms.
Proposition 5.3 (Fact 1 from [apers2022quantum]).
To prove the lower bound of $3$-colorability, the authors in [bogdanov2002lower] studied another problem called E$(3,c)$LIN-2, a problem related to deciding the satisfiability of a system of linear equations. Then the authors designed a reduction to $3$-colorability from E$(3,c)$LIN-2, which finally proves the linear query complexity lower bound for testing $3$-colorability. We will also follow a similar approach here. Let us first formally define the problem of E$(3,c)$LIN-2, where below $\mathbb{F}_{2}$ denotes the field with two elements.
@@
Lemma 5.5 (Theorem $8$ from [bogdanov2002lower]).
@@
    Each row of $A$ has exactly three non-zero entries.
@@
    Each column of $A$ has exactly $3c$ non-zero entries.
@@
    Every collection of $\delta\cdot N$ rows of $A$ is linearly independent.
Using the existence of the matrix $A$ corresponding to Lemma 5.5, the authors in [bogdanov2002lower] used Yao’s minimax lower bound technique to prove a linear lower bound for testing E$(3,c)$LIN-2. More formally, they designed a pair of hard-to-distinguish distributions $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$, such that unless $\Omega(N)$ queries are performed, no algorithm can distinguish between them. Since we will be using the same approach, we formally define $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$ below.
5.1 Testing E$(3,c)$LIN-2
@@
Given a matrix $A$ (similar to the matrix mentioned in Lemma 5.5), and query access to an unknown vector $y$, in order to distinguish if there exists another vector $x$ such that $Ax=y$, or for any vector $x$, only a constant $\varepsilon$ fraction of the constraints encoded by $A$ are satisfied for some $\varepsilon\in(0,1)$, $\Omega(N)$ queries are necessary.
Construction of hard distributions: The hard-to-distinguish distributions $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$ are as follows:
@@
    $D_{\mathrm{yes}}$: Choose a vector $z\in\{0,1\}^{N}$ uniformly at random, and set $Ay=z$.
@@
    $D_{\mathrm{no}}$: Choose two vectors $y,z\in\{0,1\}^{N}$ uniformly at random, independently of each other, and set $Ay=z$.
Now we have the following lemma describing the properties of $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$.
@@
  1. 1.
    The system of linear equations corresponding to $D_{\mathrm{yes}}$ are satisfiable.
  2. 2.
    The system of linear equations corresponding to $D_{\mathrm{no}}$ are $(1/2-\alpha)$-far from being satisfiable for every $\alpha>0$.
Proof.
The system of linear equations corresponding to $D_{\mathrm{yes}}$ are satisfiable since we can set $y=z$. On the other hand, the system of linear equations corresponding to $D_{\mathrm{no}}$ is far from being satisfiable with high probability. ∎
Now let us state the following lemma from [bogdanov2002lower], which states that the inputs drawn from $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$ will remain indistinguishable unless $\Omega(N)$ queries are performed.
Lemma 5.8 (Lemma 19 of [bogdanov2002lower]).
The key insight that is used to prove the above lemma is that applying Lemma 5.5, any $\delta\cdot N$ rows of $A$ are linearly independent, thus any subset of $\delta\cdot N$ entries of $Ay=z$ will look uniformly random. Hence $z$ is “$k$-wise independent” with $k=\delta\cdot N$. We will not formally prove the above lemma here, please refer to [bogdanov2002lower] for formal proof.
5.2 Quantum lower bound for testing E$(3,c)$LIN-2
Now we will use Proposition 5.3 to prove a linear query quantum lower bound for testing E$(3,c)$LIN-2. The idea is that unless any quantum algorithm performs $k/2$ queries, it is hard to distinguish a uniformly random vector from a $k$-wise independent vector. This implies an $\Omega(k)=\Omega(N)$ query lower bound in the quantum setting. Formally, we have a quantum variant of Lemma 5.8 stated below.
Lemma 5.9.
For every $\alpha>0$ there are constants $c$ and $\delta>0$ such that every algorithm that distinguishes satisfiable instances of E$(3,c)$LIN-2 with $N$ variables from instances that are $(1/2-\alpha)$-far from satisfiable must have quantum query complexity at least $\delta N$.
As mentioned before, using the same reduction between E$(3,c)$LIN-2 and $3$-colorability from [bogdanov2002lower, Section 5], we conclude that $\Omega(N)$ quantum queries are necessary to test $3$-colorability in the bounded degree model, thereby proving Theorem 5.1.
@@ 5.3 Other maximal hard-to-test problems
Given quantum query access to an unknown undirected $d$-bounded degree $N$-vertex graph $G$ for some integer $d$, and a parameter $\varepsilon\in(0,1)$, in order to distinguish if $G$ has a (directed) Hamiltonian path/cycle or $\varepsilon$-far from having a (directed) Hamiltonian path/cycle, $\Omega(N)$ quantum queries are necessary.
@@ 6 Conclusion
A last important point is the general framework of the dual polynomial method we made, which will hopefully initiate future research in the community.
@@ A.1 Dual polynomial for THR
Therefore, if $z_{i}=\textnormal{sgn}(\psi(x_{i}))$ and $x_{i}$ is a false $z_{i}$, it means that $z_{i}$ should be flipped to get $h(x_{i})$. Let $y_{i}\in\{-1,1\}$ denote whether we flip $z_{i}$. As $x_{i}$ is a false $z_{i}$ with probability $2\sum_{x_{i}\in D_{z_{i}}}|\psi(x_{i})|$, this is the probability with which we should flip $z_{i}$, i.e. the probability that $y_{i}=-1$.

### v2

@@ 4.2 Preparation
Notice that this problem is not a property testing problem, as the outcome is not determined based on the distance between inputs. Nevertheless, it is a valid promise problem and a special case of testing $k$-collision-freeness, which we use to prove lower bounds on the other problems of interest.
To prove a bound on the $\mathrm{Collision}$ function, we will actually relate it to the composition of two more elementary functions $g,h$, where by composition we mean $(g\circ h)(x)=g(h(x_{1}),\dots,h(x_{n}))$ (where $x=(x_{1},\dots,x_{n})$ and each $x_{i}$ is a binary vector of appropriate dimension). Let us define (i) the threshold function $\mathrm{THR}_{N}^{k}:\{-1,1\}^{N}\to\{-1,1\}$ which is $-1$ if the input bitstring contains at least $k$ many $-1$s, and it is 1 otherwise; and (2) the gap version of $\mathrm{OR}$, that is $\mathrm{GapOR}_{R}^{\gamma}:D_{\mathrm{GapOR}_{R}^{\gamma}}\subset\{-1,1\}^{R}\to\{-1,1\}$ which takes value 1 if the input is $1^{R}$, $-1$ if the input contains at least $\gamma R$ many $-1$s, and is undefined otherwise. We show that the double-promise approximate degree of $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ lower bounds the quantum query complexity of the collision problem.
@@
Notice that $(s_{1},\dots,s_{N})$ is free from $k$-collisions (ignoring collisions of the dummy character 0) if and only if $(T_{1}(s_{1}),\dots,T_{N}(s_{N}))$ is free from $k$-collisions, i.e., new $k$-collisions cannot be created by this transformation (only 2-collisions but we assume $k\geq 3$).
@@
By Proposition 4.11, instead of $\mathrm{Collision}^{k,\gamma/c}_{N,N}$ we can consider $\mathrm{dCollision}^{k,\gamma}_{N,R}$ (with the appropriate parameters) to show a lower bound. By Proposition 4.12, to relate the query complexity of $\mathrm{dCollision}^{k,\gamma}_{N,R}$ to the double-promise degree of $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ with domain further restricted to $H_{\leq N}^{NR}$. ∎
@@ 4.3 Main lower bound
We first define a construction used in order to compose dual polynomials, which was introduced in earlier line of work [ShiZ09, Lee09, Sher13].
@@
Let $N=\lceil 20(2k)^{k/2}\rceil R$ and $\phi:\{-1,1\}^{R}\to\mathbb{R}$ from Proposition 4.15. Then there exist $\psi:\{-1,1\}^{N}\to\mathbb{R}$ and $\zeta:\{-1,1\}^{NR}\to\mathbb{R}$ such that
@@
Now we focus on the composed function $f$ (and the dual composition $\phi\star\psi$), the domain is not restricted to small Hamming weight inputs anymore.
@@
Then, similarly to the first term, we can upper bound the probability. Now in the last step, we use the Chernoff bound, which introduces the constraint $\gamma<\frac{1}{4^{k-1}}$.
@@
By Lemma 4.14, there is a dual polynomial for $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ of pure high degree $\Omega(\sqrt{N^{1-1/k}}/\ln^{2}N)$ that is only supported on $H_{\leq N}^{NR}$. By Theorem 4.8 this means that the double-promise $\delta$-approximate degree of $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ with domain restricted to $H_{\leq N}^{NR}$ is $\Omega(\sqrt{N^{1-1/k}}/\ln^{2}N)$. Using Lemma 4.10 with $c=\lceil 20(2k)^{k/2}\rceil$, we obtain that the bounded-error quantum query complexity of $\mathrm{Collision}_{N,N}^{k,\gamma^{\prime}}$ is $\Omega(\sqrt{N^{1-1/k}}/\ln^{2}N)$ if $\gamma^{\prime}=\gamma/c<1/(4^{k-1}\lceil 20(2k)^{k/2}\rceil)$. This implies the same lower bound on testing $k$-collision-freeness with $\varepsilon=\gamma^{\prime}$ as $\mathrm{Collision}$ is just a more restricted version of the same problem. ∎
@@ 5 Testing $3$-colorability
In this section, we will prove that the problem of testing $3$-colorability in bounded degree graphs remains maximally hard-to-test in the quantum setting. Our lower bound proof will roughly follow the same approach as that of [bogdanov2002lower], see [bhattacharyya2022property, Section 5.6] also for a reference. Our result is stated below.
@@
In order to prove the above theorem, we will first discuss the approach to proving the classical lower bound. Then we will modify the classical proof suitably to the quantum setting. Let us start with the notion of $k$-wise independent string which will be used both in the classical and quantum lower bound proofs.
@@
5.1 Classical lower bound approach for testing $3$-colorability
As we discussed in the overview, to prove the lower bound of $3$-colorability, the authors in [bogdanov2002lower] studied another problem called E$(3,c)$LIN-2, a problem related to deciding the satisfiability of a system of linear equations. Then the authors designed a reduction to $3$-colorability from E$(3,c)$LIN-2, which finally proves the linear query complexity lower bound for testing $3$-colorability. We will also follow a similar approach here. Let us first formally define the problem of E$(3,c)$LIN-2, where below $\mathbb{F}_{2}$ denotes the field with two elements.
@@ 5.1 Classical lower bound approach for testing $3$-colorability
Lemma 5.4 ([bogdanov2002lower, Theorem 8]).
@@
    Each row of $A$ has exactly three non-zero entries;
@@
    Each column of $A$ has exactly $3c$ non-zero entries;
@@
    Every collection of $\delta N$ rows of $A$ is linearly independent.
Using the existence of the matrix $A$ corresponding to Lemma 5.4, the authors in [bogdanov2002lower] used Yao’s minimax lower bound technique to prove a linear lower bound for testing E$(3,c)$LIN-2. More formally, they designed a pair of hard-to-distinguish distributions $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$, such that unless $\Omega(N)$ queries are performed, no algorithm can distinguish between them.
@@
There exists a matrix $A\in\{0,1\}^{cN\times N}$ (similar to the matrix mentioned in Lemma 5.4) such that given a parameter $\varepsilon\in(0,1)$ and query access to $A$ and a vector $y\in\{0,1\}^{cN}$, in order to distinguish if there exists another vector $x\in\{0,1\}^{N}$ such that $Ax=y$, or for any vector $x\in\{0,1\}^{N}$, only a constant $\varepsilon$ fraction of the constraints encoded by $A$ and $y$ are satisfied, $\Omega(N)$ queries are necessary.
For completeness, we will briefly describe the proof of the above lemma in Appendix B. The authors in [bogdanov2002lower] proved the following lower bound for testing E$(3,c)$LIN-2.
Lemma 5.6 ([bogdanov2002lower, Lemma 19]).
The key insight that is used to prove the above lemma is that applying Lemma 5.4, any $\delta N$ rows of $A$ are linearly independent, thus any subset of $\delta N$ entries of $Ay=z$ will look uniformly random. Hence $z$ is “$k$-wise independent” with $k=\delta N$. We will not formally prove the above lemma here, please refer to [bogdanov2002lower] for a formal proof.
Finally, we have the reduction that maps satisfying instances of testing E$(3,c)$LIN-2 to satisfying instances of testing $3$-colorability and vice-versa.
Lemma 5.7 ([bogdanov2002lower, Section 4]).
There exists a reduction $\varphi$ that maps instances of testing E$(3,c)$LIN-2 to instances of testing 3-colorability such that the following hold:
@@
    If an input $x$ to E$(3,c)$LIN-2 is satisfiable, then $\varphi(x)$ is a $3$-colorable graph;
@@
    If an input $x$ to E$(3,c)$LIN-2 is far from being satisfiable, then $\varphi(x)$ is a graph that is far from being $3$-colorable.
5.2 Quantum lower bound for testing E$(3,c)$LIN-2 and $3$-colorability
We will first prove the quantum lower bound for testing E$(3,c)$LIN-2. Our result is stated as follows.
@@ 5.2 Quantum lower bound for testing E$(3,c)$LIN-2 and $3$-colorability
For every $\alpha>0$ there are constants $c$ and $\delta>0$ such that every algorithm that distinguishes satisfiable instances of E$(3,c)$LIN-2 with $N$ variables from instances that are $(1/2-\alpha)$-far from satisfiable must have quantum query complexity at least $\delta N/2$.
In order to prove the above lemma, we will be using the following observation from [apers2022quantum], which states that distinguishing between a uniformly random string and a $\ell$-wise independent string, for an appropriate integer $\ell$, is hard for quantum algorithms.
Proposition 5.9 (Fact 1 from [apers2022quantum]).
Now let us prove Lemma 5.8.
Proof of Lemma 5.8.
Following Lemmas 5.4 and 5.5, we know that there exists a matrix $A$ whose $\delta N$ rows are linearly independent, for which testing E$(3,c)$LIN-2 requires $\Omega(N)$ classical queries. Moreover, from Proposition 5.9, we know that any quantum algorithm that performs less than $k/2$ queries can not distinguish a uniformly random vector from a $k$-wise independent vector. Now let us set $k=\delta N$. Combining all the above, this implies that at least $\delta N/2$ quantum queries are necessary for testing E$(3,c)$LIN-2. ∎
Now we are finally ready to prove Theorem 5.1.
Proof of Theorem 5.1.
From Lemma 5.8, we know that the quantum query complexity of testing E$(3,c)$LIN-2 is $\Omega(N)$. In order to prove similar lower bound for testing $3$-colorability, we will again use a reduction approach. Given a pair of hard instances corresponding to testing E$(3,c)$LIN-2, we will apply the reduction $\psi$ mentioned in Lemma 5.7. Similar to the classical setting, $\varphi$ will map the yes instances of E$(3,c)$LIN-2 to instances of $3$-colorable graphs and vice-versa. So, the quantum query lower bound of $\Omega(N)$ carries forward from E$(3,c)$LIN-2 to $3$-colorability. Thus, we conclude that $\Omega(N)$ quantum queries are necessary to test $3$-colorability in the bounded degree model. ∎
@@ 5.3 Other maximal hard-to-test problems
Given quantum query access to an unknown undirected (directed) $d$-bounded degree $N$-vertex graph $G$ for some integer $d$, and a parameter $\varepsilon\in(0,1)$, in order to distinguish if $G$ has an undirected (directed) Hamiltonian path/cycle or $\varepsilon$-far from having an undirected (directed) Hamiltonian path/cycle, $\Omega(N)$ quantum queries are necessary.
@@ A.1 Dual polynomial for THR
Therefore, if $z_{i}=\textnormal{sgn}(\psi(x_{i}))$ and $x_{i}$ is a false $z_{i}$, it means that $z_{i}$ should be flipped to get $h(x_{i})$. Let $y_{i}\in\{-1,1\}$ denote whether we flip $z_{i}$. As $x_{i}$ is a false $z_{i}$ with probability $2\sum_{x_{i}\in D_{z_{i}}}|\psi(x_{i})|$, this is the probability with which we should flip $z_{i}$, i.e., the probability that $y_{i}=-1$.
@@
Appendix B Proofs from Section 5: hardness of testing E$(3,c)$LIN-2
In this section, we will present a sketch of the proof of the lower bound of testing E$(3,c)$LIN-2. Our main result is stated as follows.
See 5.5
Proof sketch.
As we mentioned, this proof follows Yao’s minimax lower bound technique. A pair of hard distributions $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$ are constructed, such that, unless $\Omega(N)$ queries are performed, no algorithm can distinguish between them.
Let us consider the matrix $A\in\{0,1\}^{cN\times N}$ as mentioned in Lemma 5.4. Based on the matrix $A$, the hard-to-distinguish distributions $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$ are as follows:
  1. 1.
    $D_{\mathrm{yes}}$: Choose a vector $z\in\{0,1\}^{N}$ uniformly at random from $\{0,1\}^{N}$, and set the vector $y\in\{0,1\}^{cN}$ as $y=Az$. Then the system of linear equations is $Ax=y$.
  2. 2.
    $D_{\mathrm{no}}$: Choose the vector $y\in\{0,1\}^{cN}$ uniformly at random from $\{0,1\}^{cN}$, and set the system of linear equations $Ax=y$.
Now we have the following claim describing the properties of $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$.
Claim B.1.
  1. (i)
    The system of linear equations corresponding to $D_{\mathrm{yes}}$ is satisfiable.
  2. (ii)
    With probability at least $2/3$, the system of linear equations corresponding to $D_{\mathrm{no}}$ is $(1/2-\alpha)$-far from being satisfiable for every $\alpha>0$.
Note that the system of linear equations in $D_{\mathrm{yes}}$ is satisfiable by setting $x=z$. On the other hand, for the system of linear equations corresponding to $D_{\mathrm{no}}$, vector $y$ is uniformly random. Thus, with high probability, vector $Az-y$ has large Hamming weight for any $z\in\{0,1\}^{N}$, and therefore the system of linear equations $Ax=y$ is far from being satisfiable. The formal proof is in [bogdanov2002lower, Lemma 18]. ∎